# Mustard: Mastering Uniform Synthesis of Theorem and Proof Data

### v2

@@ 4.2 Human Evaluation
Inspection Dimension
Requirement
Valid Invalid $p$-value
(D1) IS Correctness
Whether the informal statement is factually correct.
93.50 83.50 0.00167
(D2) IS Relevance
Whether the informal statement is relevant to each seed concept.
87.50 92.50 0.09604
(D3) RT Classification
Whether the informal statement is of the required question type.
67.00 68.50 0.74903
(D4) IP Correctness
Whether the informal proof correctly solves the informal statement.
88.50 73.50 0.00012
(D5) IS-FS Alignment
Whether the informal statement and the formal statement describe the same problem and are aligned with each other.
74.00 66.50 0.10138
(D6) IP-FP Alignment
Whether the informal proof and the formal proof describe the same solution and have aligned proof steps.
72.00 54.00 0.00018
@@
MODEL
Zero (G)
Few (G)
Zero (M) Few (M)
MODEL
Zero (G)
Few (G)
Zero (M) Few (M)
@@
GPT2-large
3.4
5.1
0.6 1.0
GPT2-large > gt
14.6
17.4
4.6 6.8
Llama 2-7B
7.2
12.8
2.0 2.6
Llama 2-7B > gt
24.5
28.2
10.4 12.6
@@
GPT2-large > tt
4.2
6.8
1.4
2.4
GPT2-large > tt > gt
16.1
18.9
4.8
8.0
GPT2-large > in
3.9
6.4
1.2
2.0
GPT2-large > in > gt
15.4
17.7
4.4
7.6
GPT2-large > ra
4.1
6.7
1.4
2.2
GPT2-large > ra > gt
15.7
18.5
4.8
7.8
GPT2-large > va
4.6 (+12.20%)
7.0 (+4.48%)
1.8 (+28.57%)
2.8 (+27.27%)
GPT2-large > va > gt
16.5 (+5.10%)
20.1 (+8.65%)
5.6 (+16.67%)
8.4 (+7.69%)
Llama 2-7B > tt
9.6
16.0
3.2
3.8
Llama 2-7B > tt > gt
27.4
31.5
13.0
14.4
Llama 2-7B > in
9.1
14.9
2.4
3.2
Llama 2-7B > in > gt
26.9
30.3
12.4
13.8
Llama 2-7B > ra
9.5
15.4
3.0
3.6
Llama 2-7B > ra > gt
27.1
30.7
12.6
14.2
Llama 2-7B > va
10.3 (+8.42%)
16.9 (+9.74%)
3.2 (+6.67%)
4.2 (+16.67%)
Llama 2-7B > va > gt
27.9 (+2.95%)
32.5 (+5.86%)
13.8 (+9.52%)
15.0 (+5.63%)
@@ 4.3 Data Quality by Downstream Application
MODEL
mathlib
miniF2F
test MODEL
mathlib
miniF2F
test
@@
GPT2-large
0.0
0.0
0.0
GPT2-large > mt
5.6
2.9
8.6
Llama 2-7B
0.0
0.0
0.0
Llama 2-7B > mt
14.3
7.0
10.8
@@
GPT2-large > in
2.0
0.0
6.0
GPT2-large > inmt
5.9
2.0
8.2
GPT2-large > ra
3.0
1.2
7.0
GPT2-large > ramt
6.6
2.9
9.6
GPT2-large > va
3.7 (+23.33%)
1.6 (+33.33%)
8.3 (+18.57%)
GPT2-large > vamt
7.4 (+12.12%)
3.7 (+27.59%)
10.6 (+10.42%)
@@
Llama 2-7B > in
5.8
1.2
8.6
Llama 2-7B > inmt
11.6
5.7
12.6
Llama 2-7B > ra
7.5
2.5
10.4
Llama 2-7B > ramt
14.7
6.6
13.2
Llama 2-7B > va
8.7 (+16.00%)
2.9 (+16.00%)
12.2 (+17.31%)
Llama 2-7B > vamt
15.7 (+6.80%)
7.8 (+18.18%)
14.4 (+18.18%)
@@
[Uncaptioned image] Figure 3: Model performances fine-tuned with different data scales.
@@ 4.5 Pass Rate
All (GPT-4)
Step (GPT-4)
All (GPT-3.5)
All (GPT-4)
Step (GPT-4)
All (GPT-3.5)
#correct=0 1 ($\Delta$) 2 ($\Delta$)
0
0
0 1 ($\Delta$) 2 ($\Delta$)
0
0
k=1
elem
26.0 48.0 (+22.0) 55.9 (+33.9)
25.5
15.1
38.0 59.7 (+21.7) 67.0 (+45.3)
40.1
22.2
midd
16.4 31.8 (+15.4) 39.6 (+24.2)
17.0
4.3
22.9 39.7 (+16.8) 47.4 (+30.6)
28.4
7.0
high
6.8 14.4 (+7.6) 17.2 (+9.6)
6.6
1.9
6.7 16.8 (+10.1) 21.9 (+11.8)
8.1
3.4
higher
2.1 5.6 (+3.5) 9.8 (+6.3)
3.0
0.7
3.6 10.9 (+7.3) 16.0 (+8.7)
4.3
2.7
k=2
elem
24.1 42.3 (+18.2) 52.3 (+34.1)
23.2
12.8
32.2 49.5 (+17.3) 58.2 (+40.9)
31.9
22.1
midd
14.0 25.2 (+11.2) 34.1 (+22.9)
15.0
5.3
16.9 27.3 (+10.4) 34.5 (+24.1)
17.3
5.7
high
3.8 8.3 (+4.5) 12.1 (+7.6)
5.4
2.2
5.7 11.0 (+5.3) 16.2 (+10.9)
4.5
2.5
higher
1.1 3.3 (+2.2) 5.0 (+2.8)
2.6
1.3
2.6 7.0 (+4.4) 10.5 (+6.1)
2.9
2.1
@@ Appendix B Mathematical Concepts
Domains
Domains
Domains
Domains
@@
1st grade
“Place value”, “Addition and subtraction”, “Measurement, data, and geometry”,
2nd grade
“Add and subtract within 20”, “Place value”, “Add and subtract within 100”, “Add and subtract within 1,000”, “Money and time”, “Measurement”, “Data”, “Geometry”,
3rd grade
“Intro to multiplication”, “1-digit multiplication”, “Addition, subtraction, and estimation”, “Intro to division”, “Understand fractions”, “Equivalent fractions and comparing fractions”, “More with multiplication and division”, “Arithmetic patterns and problem solving”, “Quadrilaterals”, “Area”, “Perimeter”, “Time”, “Measurement”, “Represent and interpret data”,
4th grade
“Place value”, “Addition, subtraction, and estimation”, “Multiply by 1-digit numbers”, “Multiply by 2-digit numbers”, “Division”, “Factors, multiples and patterns”, “Equivalent fractions and comparing fractions”, “Add and subtract fractions”, “Multiply fractions”, “Understand decimals”, “Plane figures”, “Measuring angles”, “Area and perimeter”, “Units of measurement”,
5th grade
“Decimal place value”, “Add decimals”, “Subtract decimals”, “Add and subtract fractions”, “Multi-digit multiplication and division”, “Multiply fractions”, “Divide fractions”, “Multiply decimals”, “Divide decimals”, “Powers of ten”, “Volume”, “Coordinate plane”, “Algebraic thinking”, “Converting units of measure”, “Line plots”, “Properties of shapes”,
6th grade
“Ratios”, “Arithmetic with rational numbers”, “Rates and percentages”, “Exponents and order of operations”, “Negative numbers”, “Variables & expressions”, “Equations & inequalities”, “Plane figures”,
@@ C.1 Prompt Template for Proof Filtering
You are a math expert. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please first write the question part regardless of the other parts. You must write the following format, filling in the ‘‘# Problem: ’’ section, and leaving the other two sections empty. # Problem: ... # Informal proof: ... # Formal proof in Lean 3: ...
You are a math expert. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please then write the corresponding solution in natural language (indicated by ‘‘Informal proof:’’) given the ‘‘# Problem: ’’, filling in the ‘‘# Informal proof: ’’ section, and leaving the other section empty. # Problem: <generated problem> # Informal proof: ... # Formal proof in Lean 3: ...
You are a master in Lean. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please translate the ‘‘# Informal proof:’’ section into Lean 3 and fill in the ‘‘# Formal proof in Lean 3: ’’ section. # Problem: <generated problem> # Informal proof: <generated informal proof> # Formal proof in Lean 3: ...
@@ Appendix E Case Study
def ABCD_is_square : $\forall$ (AB BC CD DA : $\mathbb{R}$), AB = 5 $\land$ BC = 5 $\land$ CD = 5 $\land$ DA = 5 normal-$\to$ is_square_quad AB BC CD DA :=
@@
[…]
@@
def f := $\lambda$(t : $\mathbb{R}$), ${}^{2}$ + 3*t + 1
lemma f_value_at_90 : f 90 = ${}^{2}$ + 3*(90) + 1 :=
@@
($\exists$ x, diff_eq p q y1 x $\land$ diff_eq p q y2 x) normal-$\to$ ($\forall$ x, diff_eq p q y x) normal-$\to$
@@
lemma matrix_times_differentiable_function : differentiable_at $\mathbb{R}$ ($\lambda$ x, A normal-$\cdot$ ($\lambda$ i, f x) i) $x_{0}$ :=
@@
lemma derivative_of_matrix_times_function : deriv ($\lambda$ x, A normal-$\cdot$ ($\lambda$ i, f x) i) $x_{0}$ = A normal-$\cdot$ ($\lambda$ i, deriv f $x_{0}$) i :=
@@ Appendix H Data Contamination Check
[…]
@@
[…]
@@
[…]

### v3

@@ 4.2 Human Evaluation
Inspection Dimension Requirement Valid Invalid $p$-value
(D1) IS Correctness Whether the informal statement is factually correct. 93.50 83.50 0.00167
(D2) IS Relevance Whether the informal statement is relevant to each seed concept. 87.50 92.50 0.09604
(D3) RT Classification Whether the informal statement is of the required question type. 67.00 68.50 0.74903
(D4) IP Correctness Whether the informal proof correctly solves the informal statement. 88.50 73.50 0.00012
(D5) IS-FS Alignment Whether the informal statement and the formal statement describe the same problem and are aligned with each other. 74.00 66.50 0.10138
(D6) IP-FP Alignment Whether the informal proof and the formal proof describe the same solution and have aligned proof steps. 72.00 54.00 0.00018
@@
MODEL Zero (G) Few (G) Zero (M) Few (M) MODEL Zero (G) Few (G) Zero (M) Few (M)
@@
GPT2-large 3.4 5.1 0.6 1.0 GPT2-large > gt 14.6 17.4 4.6 6.8
Llama 2-7B 7.2 12.8 2.0 2.6 Llama 2-7B > gt 24.5 28.2 10.4 12.6
@@
GPT2-large > tt 4.2 6.8 1.4 2.4 GPT2-large > tt > gt 16.1 18.9 4.8 8.0
GPT2-large > in 3.9 6.4 1.2 2.0 GPT2-large > in > gt 15.4 17.7 4.4 7.6
GPT2-large > ra 4.1 6.7 1.4 2.2 GPT2-large > ra > gt 15.7 18.5 4.8 7.8
GPT2-large > va 4.6 (+12.20%) 7.0 (+4.48%) 1.8 (+28.57%) 2.8 (+27.27%) GPT2-large > va > gt 16.5 (+5.10%) 20.1 (+8.65%) 5.6 (+16.67%) 8.4 (+7.69%)
Llama 2-7B > tt 9.6 16.0 3.2 3.8 Llama 2-7B > tt > gt 27.4 31.5 13.0 14.4
Llama 2-7B > in 9.1 14.9 2.4 3.2 Llama 2-7B > in > gt 26.9 30.3 12.4 13.8
Llama 2-7B > ra 9.5 15.4 3.0 3.6 Llama 2-7B > ra > gt 27.1 30.7 12.6 14.2
Llama 2-7B > va 10.3 (+8.42%) 16.9 (+9.74%) 3.2 (+6.67%) 4.2 (+16.67%) Llama 2-7B > va > gt 27.9 (+2.95%) 32.5 (+5.86%) 13.8 (+9.52%) 15.0 (+5.63%)
@@ 4.3 Data Quality by Downstream Application
MODEL mathlib miniF2F test MODEL mathlib miniF2F test
@@
GPT2-large 0.0 0.0 0.0 GPT2-large > mt 5.6 2.9 8.6
Llama 2-7B 0.0 0.0 0.0 Llama 2-7B > mt 14.3 7.0 10.8
@@
GPT2-large > in 2.0 0.0 6.0 GPT2-large > inmt 5.9 2.0 8.2
GPT2-large > ra 3.0 1.2 7.0 GPT2-large > ramt 6.6 2.9 9.6
GPT2-large > va 3.7 (+23.33%) 1.6 (+33.33%) 8.3 (+18.57%) GPT2-large > vamt 7.4 (+12.12%) 3.7 (+27.59%) 10.6 (+10.42%)
@@
Llama 2-7B > in 5.8 1.2 8.6 Llama 2-7B > inmt 11.6 5.7 12.6
Llama 2-7B > ra 7.5 2.5 10.4 Llama 2-7B > ramt 14.7 6.6 13.2
Llama 2-7B > va 8.7 (+16.00%) 2.9 (+16.00%) 12.2 (+17.31%) Llama 2-7B > vamt 15.7 (+6.80%) 7.8 (+18.18%) 14.4 (+18.18%)
@@
Figure 3: Model performances fine-tuned with different data scales.
@@ 4.5 Pass Rate
All (GPT-4) Step (GPT-4) All (GPT-3.5) All (GPT-4) Step (GPT-4) All (GPT-3.5)
#correct=0 1 ($\Delta$) 2 ($\Delta$) 0 0 0 1 ($\Delta$) 2 ($\Delta$) 0 0
k=1 elem 26.0 48.0 (+22.0) 55.9 (+33.9) 25.5 15.1 38.0 59.7 (+21.7) 67.0 (+45.3) 40.1 22.2
midd 16.4 31.8 (+15.4) 39.6 (+24.2) 17.0 4.3 22.9 39.7 (+16.8) 47.4 (+30.6) 28.4 7.0
high 6.8 14.4 (+7.6) 17.2 (+9.6) 6.6 1.9 6.7 16.8 (+10.1) 21.9 (+11.8) 8.1 3.4
higher 2.1 5.6 (+3.5) 9.8 (+6.3) 3.0 0.7 3.6 10.9 (+7.3) 16.0 (+8.7) 4.3 2.7
k=2 elem 24.1 42.3 (+18.2) 52.3 (+34.1) 23.2 12.8 32.2 49.5 (+17.3) 58.2 (+40.9) 31.9 22.1
midd 14.0 25.2 (+11.2) 34.1 (+22.9) 15.0 5.3 16.9 27.3 (+10.4) 34.5 (+24.1) 17.3 5.7
high 3.8 8.3 (+4.5) 12.1 (+7.6) 5.4 2.2 5.7 11.0 (+5.3) 16.2 (+10.9) 4.5 2.5
higher 1.1 3.3 (+2.2) 5.0 (+2.8) 2.6 1.3 2.6 7.0 (+4.4) 10.5 (+6.1) 2.9 2.1
@@ Appendix B Mathematical Concepts
Domains # Domains # Domains # Domains #
@@
1st grade “Place value”, “Addition and subtraction”, “Measurement, data, and geometry”,
2nd grade “Add and subtract within 20”, “Place value”, “Add and subtract within 100”, “Add and subtract within 1,000”, “Money and time”, “Measurement”, “Data”, “Geometry”,
3rd grade “Intro to multiplication”, “1-digit multiplication”, “Addition, subtraction, and estimation”, “Intro to division”, “Understand fractions”, “Equivalent fractions and comparing fractions”, “More with multiplication and division”, “Arithmetic patterns and problem solving”, “Quadrilaterals”, “Area”, “Perimeter”, “Time”, “Measurement”, “Represent and interpret data”,
4th grade “Place value”, “Addition, subtraction, and estimation”, “Multiply by 1-digit numbers”, “Multiply by 2-digit numbers”, “Division”, “Factors, multiples and patterns”, “Equivalent fractions and comparing fractions”, “Add and subtract fractions”, “Multiply fractions”, “Understand decimals”, “Plane figures”, “Measuring angles”, “Area and perimeter”, “Units of measurement”,
5th grade “Decimal place value”, “Add decimals”, “Subtract decimals”, “Add and subtract fractions”, “Multi-digit multiplication and division”, “Multiply fractions”, “Divide fractions”, “Multiply decimals”, “Divide decimals”, “Powers of ten”, “Volume”, “Coordinate plane”, “Algebraic thinking”, “Converting units of measure”, “Line plots”, “Properties of shapes”,
6th grade “Ratios”, “Arithmetic with rational numbers”, “Rates and percentages”, “Exponents and order of operations”, “Negative numbers”, “Variables & expressions”, “Equations & inequalities”, “Plane figures”,
@@ C.1 Prompt Template for Proof Filtering
(T1) You are a math expert. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please first write the question part regardless of the other parts. You must write the following format, filling in the ‘‘# Problem: ’’ section, and leaving the other two sections empty. # Problem: ... # Informal proof: ... # Formal proof in Lean 3: ...
(T2) You are a math expert. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please then write the corresponding solution in natural language (indicated by ‘‘Informal proof:’’) given the ‘‘# Problem: ’’, filling in the ‘‘# Informal proof: ’’ section, and leaving the other section empty. # Problem: <generated problem> # Informal proof: ... # Formal proof in Lean 3: ...
(T3) You are a master in Lean. Now please come up with a math problem according to the following requirements. The math problem should contain a question part (indicated by ‘‘Problem: ’’), a corresponding solution in natural language (indicated by ‘‘Informal proof:’’), and a translated formal solution in Lean 3 (indicated by ‘‘Formal proof in Lean 3:’’). Please note that the informal proof and the formal proof need to be identical. Please create a [QUESTION TYPE] in the level of [EDUCATIONAL LEVEL] based on the following knowledge point(s): [CONCEPT] in [DOMAIN]; [CONCEPT] in [DOMAIN]. Please translate the ‘‘# Informal proof:’’ section into Lean 3 and fill in the ‘‘# Formal proof in Lean 3: ’’ section. # Problem: <generated problem> # Informal proof: <generated informal proof> # Formal proof in Lean 3: ...
@@ Appendix E Case Study
def ABCD_is_square : $\forall$ (AB BC CD DA : $\mathbb{R}$), AB = 5 $\land$ BC = 5 $\land$ CD = 5 $\land$ DA = 5 $\to$ is_square_quad AB BC CD DA :=
@@
cases h with h1 h2,
cases h2 with h3 h4,
cases h4 with h5 h6,
@@
def f := $\lambda$(t : $\mathbb{R}$), 2*t2 + 3*t + 1
lemma f_value_at_90 : f 90 = 2*(90)2 + 3*(90) + 1 :=
@@
($\exists$ x, diff_eq p q y1 x $\land$ diff_eq p q y2 x) $\to$ ($\forall$ x, diff_eq p q y x) $\to$
@@
lemma matrix_times_differentiable_function : differentiable_at $\mathbb{R}$ ($\lambda$ x, A $\cdot$ ($\lambda$ i, f x) i) $x_{0}$ :=
@@
lemma derivative_of_matrix_times_function : deriv ($\lambda$ x, A $\cdot$ ($\lambda$ i, f x) i) $x_{0}$ = A $\cdot$ ($\lambda$ i, deriv f $x_{0}$) i :=
@@ Appendix H Data Contamination Check
GOAL   $\alpha\text{: Type u}$,  $\beta\text{: Type v}$ ,  $\gamma\text{ : Type w}$,  f : $\text{f : }\alpha\to\beta\to\gamma$, $\text{l}_{1}$ : list $\alpha\vdash\forall(\text{l}_{2}:\text{list}\ \beta)$, (list.map2 f $\text{l}_{1}$ $\text{l}_{2}$).length = linear_order.min $\text{l}_{1}$.length $\text{l}_{2}$.length
PROOFSTEP induction l1; intro l2; cases l2; simp [*, add_one, min_succ_succ, nat.zero_min, nat.min_zero]
@@
Given two vectors u = (u1,u2) and v = (v1,v2) in R2, prove that the operation of vector addition, defined as u + v = (u1+v1, u2+v2), is commutative. That is, prove that for all u and v in R2, u + v = v + u.
@@
Given two vectors u = (u1,u2) and v = (v1,v2) in R2, prove that the operation of vector addition, defined as u + v = (u1+v1, u2+v2), is commutative. That is, prove that for all u and v in R2, u + v = v + u.